OCTET 1 OCTET 2 OCTET 3 OCTET 3 COMMENTS ADDRESS MASK SUBNET NUMBER FIRST ADDRESS BROADCAST 10 0 255 248 LAST ADDRESS
Answer #1
OCTET 1 OCTET 2 OCTET 3 OCTET 3 COMMENTS
ADDRESS 10 1 1 1 0000 1010.0000 0001.0000 0001.0000 0001
MASK 255 248 0 0 1111 1111. 1111 1000.0000 0000.0000 0000
SUBNET NUMBER 10 0 0 0

Apply 1+0 or 0+1 or 0+0= 0, 1+1=1

0000 1010.0000 0000.0000 0000.0000 0000

FIRST ADDRESS 10 0 0 0

Same as subnet number

0000 1010.0000 0000.0000 0000.0000 0000

BROADCAST 10 7 255 255

Mask is

1111 1111. 1111 1000.0000 0000.0000 0000

Address is

0000 1010.0000 0001.0000 0001.0000 0001

So, whereever there is a zero in the mask, keep it

as 1 in place of that and we will get the broadcast address.

Where there is 1 in the mask, keep the broadcast address same as address.

0000 1010.0000 0111.1111 1111. 1111 1111
LAST ADDRESS 0 1 1 1

Mask is

1111 1111. 1111 1000.0000 0000.0000 0000

Complement of mask is

0000 0000.0000 0111.1111 1111.1111 1111

Address is

0000 1010.0000 0001.0000 0001.0000 0001

So, add the complement and address using 1+0 or 0+1 or 0+0= 0, 1+1=1 and you will get the below

0000 0000.0000 0001.0000 0001.0000 0001